# SBoRA: Low-Rank Adaptation with Regional Weight Updates

### v2

@@ 5.2 Evaluating SBoRA on Arithmetic Reasoning
We also evaluated the effectiveness of SBoRA on the Arithmetic Reasoning task using the LLaMA-7B and LLaMA3-8B models. The fine-tuning and evaluation datasets are constructed following the dataset settings in LLM-Adapters [11]. We performed tests with ranks 32 and 64. As the training set for arithmetic reasoning is smaller (10k) compared to the commonsense reasoning training set, we increased the number of training epochs to 3 for better convergence and set the learning rate to 3e-4.
Table 4 presents the performance of PEFT methods, including GPT-3.5. On average, for LLaMA-7B, GPT-3.5 (175B) outperforms adapter-based PEFT LLMs in accuracy. However, for simpler math reasoning tasks like MultiArith, adapter-based methods outperform GPT-3.5, and SBoRA-FA achieves the best performance. Initially, both SBoRA-FA and SBoRA-FB used a rank of 64 for consistency in trainable parameters. SBoRA-FA achieves the best average accuracy. Increasing the rank of other baselines to 64 improves accuracy but still lags behind SBoRA-FA. Reducing SBoRA’s rank to 32 maintains comparable performance to rank-32 baselines. Notably, for the challenging task AQuA, SBoRA-FA with rank 32 excels. Moving to LLaMA3-8B, SBoRA-FA with rank 32 achieves the best performance, followed by SBoRA-FA with rank 64.
SBoRA-FA/FB achieves comparable or superior results to other baseline PEFT methods with half the trainable parameters for the same rank. Arithmetic reasoning tasks are particularly well-suited for SBoRA due to the specialized knowledge required in mathematics. SBoRA’s regional weight update property makes it more suitable for higher ranks like 64, impacting more pre-trained weight regions.
Table 4: Comparison of LLaMA-7B and LLaMA3-8B with different PEFT methods, and evaluating on the arithmetic reasoning task. The first row list out the results of GPT-3.5 for reference. We report the accuracy results for each of the sub-tasks as well as the average accuracy, higher is better for all metrics. The number of trainable parameters (TP) can be found in each row.
@@ 5.3 QSBoRA Evaluation on MMLU
We explored the effectiveness of SBoRAs in the memory-efficient QLoRA framework [4]. QLoRA utilizes techniques including 4-bit NormalFloat (NF4) double quantization, gradient checkpointing, and paged optimizer to reduce memory usage. Following the experiments in QLoRA, we fine-tuned the NF4 quantized versions of LLaMA-7B/13B and LLaMA3-8B using the Alpaca [24] and sampled FLAN v2 [28] datasets. Model performance was evaluated on the MMLU benchmark [8] by comparing the mean 5-shot MMLU test accuracy.
To ensure fairness, we used the same hyperparameter settings for all experiments. For the Alpaca dataset, we applied a batch size (bs) of 16 and 10k training steps. To maintain a similar number of training epochs with Alpaca, for FLAN v2, we used batch size 8 and 20k training steps. In the QLoRA framework, we compared SBoRA with LoRA and DoRA, using rank 64 for all methods. Adapters were added to linear layers in both self-attention and MLP modules.
The results shown in Table 5 reveal that within LLaMA-7B, both QSBoRA-FA/FB achieve comparable accuracy results to QLoRA. In the case of LLaMA-13B and LLaMA3-8B, both QSBoRA-FA/FB outperform QLoRA and QDoRA in terms of accuracy, indicating its compatibility with larger-scale models.
Table 5: Finetune LLaMA-7B/13B and LLaMA3-8B on Alpaca and Flan v2. We measured the performance using MMLU benchmark and report the 5-shot average accuracy. The training settings and the number of trainable parameters (TP) are included.
@@
PEFT TP Alpaca Flanv2 TP Alpaca Flanv2 TP Alpaca Flanv2
QLoRA 80.0M 37.9 44.4 125.2M 45.4 46.7 83.9M 51.9 49.5
@@ 6 Conclusion
We introduce SBoRA, a novel approach for parameter-efficient fine-tuning of LLMs. Our proposed SBoRA-FA and SBoRA-FB methods utilize the low-rank projection-down and projection-up matrices constructed from standard basis vectors, leveraging one-dimensional index representations and sample-based multiplication. This innovative approach reduces memory and computational requirements while enhancing learning capacity, outperforming LoRA and matching the performance of DoRA. Empirical results demonstrate that SBoRA-FA is the preferred method for single-task LLM fine-tuning, conserving activation memory while achieving superior performance.
The SBoRA approach enables regional weight updates, preserving most of the pre-trained model weights while efficiently adapting to new tasks. This localized learning process draws parallels with the modular organization of the brain, where distinct cognitive functions are localized to specific brain regions. This analogy highlights the potential of SBoRA to inspire AI architectures that mimic the efficiency and adaptability of biological neural systems.
SBoRA holds immense potential for further development, particularly in multi-task training. The introduction of Multi-SBoRA would create a powerful framework for efficient adaptation to multiple tasks. Each task would undergo independent, non-overlapping weight fine-tuning with SBoRA, allowing the integration of task-specific knowledge while minimizing interference and maximizing the model’s capacity to leverage shared information. This approach enables the maintenance of distinct capabilities for each task within a single model, paving the way for more efficient and effective AI systems

### v3

@@ 5.2 Evaluating SBoRA on Arithmetic Reasoning
We also evaluated the effectiveness of SBoRA on arithmetic reasoning tasks using the LLaMA-7B and LLaMA3-8B models. We constructed the fine-tuning and evaluation datasets following the dataset settings in LLM-Adapters [11]. We performed tests with ranks 32 and 64. Since the training set for arithmetic reasoning is smaller (10k) compared to the commonsense reasoning training set, we increased the number of training epochs to 3 for better convergence and set the learning rate to 3e-4.
Table 4 presents the performance of PEFT methods, including GPT-3.5. On average, for LLaMA-7B, GPT-3.5 (175B) outperforms adapter-based PEFT LLMs in accuracy. However, for simpler math reasoning tasks like MultiArith, adapter-based methods outperform GPT-3.5, and SBoRA-FA achieves the best performance. Initially, SBoRA-FA/FB, OA, and LoRA-FA used a rank of 64 for consistency in trainable parameters. SBoRA-FA achieves the best average accuracy. Increasing the rank of other baselines to 64 improves accuracy but still lags behind SBoRA-FA. We then reduced the rank of SBoRA, OA, and LoRA-FA to 32, it can be found that SBoRA maintains comparable performance to rank 32 baselines. Notably, for the challenging task AQuA, SBoRA-FA with rank 32 excels. Moving to LLaMA3-8B, SBoRA-FA with rank 32 achieves the best performance, followed by SBoRA-FA with rank 64. Notably, SBoRA-FA consistently outperforms OA and LoRA-FA with both rank 32 and 64.
Table 4: Comparative analysis of LLaMA-7B/LLaMA3-8B with different Parameter-Efficient Fine-Tuning methods, evaluating on the arithmetic reasoning tasks. Results of GPT-3.5 are provided in the first row for reference. We report accuracy (%) for eight sub-tasks and average accuracy (%), with higher values indicating better performance. Column headers denote TP for trainable parameters and $r$ for rank.
@@
SBoRA-FA/FB achieves comparable or superior results to other baseline PEFT methods with half or same the trainable parameters for the same rank. Arithmetic reasoning tasks are particularly well-suited for SBoRA due to the specialized knowledge required in mathematics. SBoRA’s regional weight update property makes it more suitable for higher ranks like 64, impacting more pre-trained weight regions.
Based on experimental observations, SBoRA has been shown to significantly reduce training time compared to DoRA. To further illustrate the computational and memory efficiency advantages of SBoRA, we conducted a detailed analysis of the time cost and GPU memory usage during the training process for arithmetic reasoning tasks. The results are visually represented in the accompanying diagrams, which were performed on an NVIDIA RTX 4090.
We report the training time and GPU memory allocation for both LLaMA-7B and LLaMA3-8B models, comparing SBoRA with LoRA, LoRA-FA, and DoRA using ranks of 32 and 64 for each method. Figures 3 and 4 present the results for LLaMA-7B and LLaMA3-8B, respectively. The diagrams clearly demonstrate that, at equivalent ranks, SBoRA exhibits reduced training time and memory usage compared to both LoRA and DoRA during the training process. Notably, SBoRA-FA and SBoRA-FB achieve a significant reduction in training time, more than halving the duration required by DoRA.
Figure 3: GPU usage and training time for LLaMA-7B on arithmetic reasoning tasks. Results for rank 64 (left) and 32 (right) are displayed. Y-axis: GPU usage; X-axis: training time. Total training time is labeled for each method.
Figure 4: GPU usage and training time for LLaMA3-8B on arithmetic reasoning tasks. Results for rank 64 (left) and 32 (right) are displayed. Y-axis: GPU usage; X-axis: training time. Total training time is labeled for each method.
@@ 5.3 QSBoRA Evaluation on MMLU
In this section, we explored the effectiveness of SBoRAs in the memory-efficient QLoRA framework [4]. QLoRA utilizes techniques such as 4-bit NormalFloat (NF4) double quantization, gradient checkpointing, and paged optimizer to reduce memory usage. Following the experiments in QLoRA, we fine-tuned the NF4 quantized versions of LLaMA-7B/13B and LLaMA3-8B using the Alpaca [24] and sampled FLAN v2 [28] datasets. We evaluated model performance on the MMLU benchmark [8] by comparing the mean 5-shot MMLU test accuracy.
Table 5: Finetune LLaMA-7B/13B and LLaMA3-8B on Alpaca and Flan v2. We measured the performance using MMLU benchmark and report the 5-shot average accuracy. The training sets and the number of trainable parameters (TP) are included.
@@
PEFT $r$ TP Alpaca Flanv2 TP Alpaca Flanv2 TP Alpaca Flanv2
QLoRA 64 80.0M 37.9 44.4 125.2M 45.4 46.7 83.9M 51.9 49.5
@@
QSBoRA-FA 128 87.0M 39.3 45.8 136.3M 50.0 51.7 88.1M 56.3 54.2
QSBoRA-FB 72.9M 39.0 45.7 114.0M 49.5 50.5 79.7M 54.5 52.4
To ensure fairness, we used the same hyperparameter settings for all experiments. For the Alpaca dataset, we used a batch size (bs) of 16 and 10k training steps. For FLAN v2, we used a batch size 8 and 20k training steps. In the QLoRA framework, we initially used a rank of 64 for all methods, including SBoRA-FA/FB, LoRA, DoRA, OA, and LoRA-FA. We then doubled the rank to 128 for SBoRA-FA/FB to maintain a similar number of trainable parameters as LoRA and DoRA with rank 64. We added adapters to linear layers in both self-attention and MLP modules.
The results in Table 5 show that within LLaMA-7B, both QSBoRA-FA/FB with rank 64 achieve comparable accuracy results to QLoRA. In the case of LLaMA-13B and LLaMA3-8B, QSBoRA-FA with rank 64 outperforms other baseline methods in terms of accuracy, indicating its compatibility with larger-scale models. Furthermore, increasing the rank to 128 leads to improved performance for both QSBoRA-FA and QSBoRA-FB in LLaMA-7B and LLaMA-13B.
@@ 6 Conclusion
We presented SBoRA, an innovative PEFT method for LLMs that boosts learning capacity while cutting down on memory and computational needs. Our experiments demonstrate that SBoRA-FA excels in both commonsense and arithmetic reasoning tasks, even when using roughly half the number of trainable parameters, and significantly reduces training time and GPU memory usage. Empirical results indicate that SBoRA-FA surpasses methods like LoRA, LoRA-FA, and OA, and matches the performance of DoRA, making it an excellent choice for single-task fine-tuning. Within the QLoRA framework, SBoRA-FA is highly compatible with quantized models and further enhances performance at higher ranks, proving to be a robust, versatile, and resource-efficient fine-tuning method.
SBoRA efficiently adapts to new tasks while preserving pre-trained weights, akin to the modular organization of the brain. This approach could inspire AI architectures that emulate biological neural systems. Future work includes developing Multi-SBoRA, which enables independent fine-tuning for each task, minimizing interference and maximizing task-specific knowledge. This advancement could lead to more efficient AI systems with distinct capabilities across multiple tasks.